# Hidden Breakthroughs in Language Model Training

### v2

@@ Abstract
Loss curves are smooth during most of model training, so visible discontinuities stand out as possible conceptual breakthroughs. Studying these breakthroughs enables a deeper understanding of learning dynamics, but only when they are properly identified. This paper argues that similar breakthroughs occur frequently throughout training but they are obscured by a loss metric that collapses all variation into a single scalar. To find these hidden transitions, we introduce POLCA, a method for decomposing changes in loss along arbitrary bases of the low-rank training subspace. We use our method to identify clusters of samples that share similar changes in loss during training, disaggregating the overall loss into that of smaller groups of conceptually similar data. We validate our method on synthetic arithmetic and natural language tasks, showing that POLCA recovers clusters that represent interpretable breakthroughs in the model’s capabilities. We demonstrate the promise of these hidden phase transitions as a tool for unsupervised interpretability.
@@ 1 Introduction
As large language models train, various internal structures develop during abrupt phase transitions. These sudden drops in loss reveal the formation of mechanisms for in-context learning [25], natural language grammar [6], hierarchical generalization [21], and many other concepts [19, 16, 27, 1]. However, the loss curve as a whole remains stubbornly smooth. As a result, these momentary conceptual breakthroughs are treated as isolated curiosities, while the majority of training behavior is considered predictable.
This work shows that in fact, a model undergoes many breakthroughs during training, most which are concealed when averaging all data into a single loss curve. Instead, we divide up the loss curve in two different ways to find hidden breakthroughs. First, we disaggregate the aggregate loss into losses on individual examples. By clustering the individual loss curves, we identify subsets of data that experience synchronized changes in loss, implying that they rely on the same conceptual breakthrough. However, any individual example might benefit from multiple breakthroughs; such an example may undergo changes synchronized with different data subsets at different times. Furthermore, distinct concepts might appear simultaneously, erroneously merging their data clusters. This means we may have to identify multiple separate breakthroughs which affect the loss curve of a single example.
@@
We show that some learned concepts can be identified by clustering exact loss, while others cannot (Section 4.2).
Using POLCA, we extend our cluster analysis to identify conceptual breakthroughs in a restricted gradient subspace that are obscured in the exact loss curves. We automatically identify specific concepts learned during breakthroughs in both synthetic (Section 4) and natural language settings (Section 5).
@@ 2 Background: How much can we learn from learning dynamics?
Various loss phase transitions have been interpreted as conceptual breakthroughs. But why expect additional interpretable breakthroughs to underlie periods of undifferentiated, gradual model improvement? Our approach is justified by the nature of the loss surface’s complexity, illlustrated by Figure 1, in which a smooth curve emerges by eliding phase transitions from each dimension.
Why expect multiple phase transitions?
A very early phase transition is to be expected early in training after a brief memorization stage [30]. In this sense, the most celebrated breakthroughs are not the first phase transitions in their training curves. Some breakthroughs even depend on earlier breakthroughs, as observed in synthetic tasks [1] and in grammar acquisition [6]. If one concept depends on another, each must appear at a different timestep, requiring multiple phase transitions.
Multiple phase transitions can also come from differences in gradient scale along different directions. Ma et al. [18] even attributed the early edge-of-stability phase transition [13, 8] to multiscale structure of the loss surface and, furthermore, noted that this multiscale structure emerges at the range where models become singular: their loss lacks a quadratic approximation in terms of model parameters, creating conditions for phase transitions under Singular Learning Theory [33, 34, 32]. They argued that this structure is the product of both nonuniform data and nonconvex objectives, respectively justifying the disaggregation and decomposition which we apply to interpret training dynamics.
@@
We track learning on training datapoints and subpopulations, rather than the whole training set, because relevant skills can be acquired at different rates [3, 7]. Individual samples thus exhibit changes in loss out of line with the monotonic average trend [36, 29]. In full-batch gradient descent, Cohen et al. [8] identified non-monotonicity arising from oscillation about the maximum Hessian eigenvector. Rosenfeld and Risteski [29] demonstrated that these oscillations occur across different axes for different samples and identified the primary cause: surprisingly human-interpretable semantic features. Even when the loss seems stable, performance can oscillate on edge cases until the model develops relevant capabilities [28, 4]. We hypothesize that oscillation represents competing skills that are relevant to different subsets of data. To test this hypothesis, and to interpret the meaning of these directions, we disaggregate the loss into clusters with similar dynamics.
@@
Our POLCA decomposition addresses what Michaud et al. [20] called polygenic scaling effects—samples which combine multiple skills and therefore exhibit breakthroughs at multiple scales. If we assume that a specific skill is enabled by movement along a particular skill’s basis vector, then the loss change attributed to movement along that vector will accelerate at the moment the skill is acquired, for every sample that requires that skill. In this manner, the sample transitions from early to late dynamics through a basis-specific loss phase transition. In other words, by monitoring changes in directions corresponding to specific skills, we support the speculation of Nanda et al. [23] that “phase transitions are everywhere.”
@@
In practice, a conceptual breakthrough might not occur in a single direction that persists throughout training. However, there is an abundance of evidence that linear bases of the low-rank [12] training subspace are conceptually meaningful. In the late stages of training, loss is convex on the line connecting a pair of checkpoints [10] if those checkpoints express similar capabilities [14] and mechanisms [17]. If a pair of high-dimensional models lack this linear connection, they still connect nonlinearly [9]; however, while parameter settings sampled from their linear connections improve broadly on the capabilities of the original models, those sampled from their nonlinear connections are less robust than the originals [14, ref Appendix D]. These observations suggest that linear decomposition should preserve meaningful conceptual features on the loss surface, and our experiments show that the resulting directions are interpretable in practice.
@@ 3 Methods
The key to our approach is the separate consideration of how each example’s datapoint loss changes throughout training. We contrast this individualized metric with aggregated loss across an entire dataset. Using the datapoint loss, we can cluster individual examples on the basis of their loss , change in loss , or magnitude of change during training.
Our objective is to decompose the loss itself into specific directions in the weight space, motivated by several considerations: First, while we have moved from an aggregated loss metric to a more granular datapoint loss metric, we are still only considering breakthroughs that are general enough to be perceived in loss curves. Second, an individual datapoint may benefit from a variety of conceptual breakthroughs, but will not be clustered on the breakthroughs individually. Finally, once we have identified a subset of the data as benefiting from a particular conceptual breakthrough, decomposing into individual weight directions allows us to locate where in the weights the breakthrough occurs and to thereby identify the mechanism involved.
Next we break this loss down by directional movement during training, allowing us to discover breakthroughs that are specific to a given direction. Our procedure follows three steps: (1) select a basis, (2) decompose the loss along that basis to highlight particular learning events, (3) cluster datapoints according to their shared learning events.
@@ 3.1 Finding the basis
To decompose the loss, we first require an interpretable orthogonal basis. We efficiently compute the eigenvectors of the Hessian matrix with CoLA [26] and use them to construct a restricted training subspace. We posit this basis to be interpretable because each basis vector expresses a high gradient covariance and therefore represents a potential decision boundary. While we select this basis for interpretability, our approach can use any arbitrary choice of basis.
@@
This basis is constructed as shown in Algorithm 1. Given intermediate training checkpoints and a number of eigenvectors to compute at each checkpoint, we seek a low rank -dimensional subspace which captures most of the movement during optimization [12]. We construct this basis iteratively, starting with : at each checkpoint , we take checkpoint weights and project their loss Hessian onto the nullspace of . From the resulting projection, we append the top eigenvectors to . We compute the eigenvectors using Hessian-vector products [11] to avoid explicitly constructing the full Hessian matrix. The resulting basis is designed to include directions of highest curvature at each checkpoint so that it will capture synchronized loss behavior throughout training. Note that the very top eigenvectors are likely to reflect local oscillation, rather than conceptually meaningful long-term movement [31], but as we continue to add to the low rank basis, we include more directions of long-term stable movement. We discard the oscillatory directions which do not provide an overall decrease in loss over the course of training according to POLCA.
3.2 Decomposing the loss with Projection Oriented Loss Change Allocation (POLCA)
@@
To decompose the loss along our basis, we propose a modified version of Loss Change Allocation (LCA) [15]. LCA is a tool for analyzing changes in aggregated loss on dataset between two checkpoints. The output of LCA is the empirical loss change between a pair of checkpoints which can be attributed to the motion of each individual weight unit. Given two consecutive checkpoints with parameters and , LCA reformulates the change in loss as its first-order Taylor approximation, a sum of the loss changes attributed to the movement of each individual model parameter :
@@
The POLCA decomposition differs from LCA in three key ways. First, we do not restrict each direction to correspond to a single unit , instead permitting an arbitrary orthonormal basis vector to replace the axis-aligned basis vectors in LCA; we project onto this basis vector using the dot product . Second, we are interested in changes in the loss on each individual example , not the entire dataset . These first two modifications provide the following reformulation of LCA.
@@
The third key difference is that we must use a second-order approximation because this basis is constructed explicitly from the Hessian eigenvectors. To understand why this choice of basis requires a second-order approximation, recall that each basis vector is an eigenvector of the Hessian matrix at some timestep , where is chosen because it has the largest eigenvalue over the whole dataset. If we assume that the top eigenvectors of the aggregate Hessian maintain high curvature at other points in training and on individual datapoints, then the scaling factor in the second-order Taylor term will be very large even at the datapoint level. Limiting the approximation to only the first order term gives poor guarantees on error, as the second-order term could be expected to dominate. We find that empirically, the difference between the first and second-order values is small (Appendix F), but compute the second-order approximation to guarantee a better estimate.
@@
Exact computation of the second-order term would be intractable, requiring computation of the top eigenvalues/vectors for each individual datapoint . Instead, we can approximate it by substituting the true eigenvalue, denoted , with the curvature of the individual loss in the direction , i.e. . If the aggregate Hessian eigenvector is close to the span of the top eigenvectors of the datapoint-specific Hessian for , this provides a reasonable estimate while reducing calculation to a single Hessian-vector product per eigenvector. We therefore approximate the basis projection of the datapoint Hessian as derived in Appendix A.
@@
Equipped with this second-order approximation of the datapoint Hessian’s projection onto our basis, we account for the high curvature and possible domination by the higher order term by modifying Equation 4 into the second-order Taylor expansion using the approximation from Equation 7.
@@ 3.3 Clustering the loss
POLCA, above, provides curves that show how a decomposed loss changes with respect to each training example. We assume that if several examples show similarly timed loss changes in the same direction, they likely rely on the same conceptual breakthroughs or learning events; therefore, they are likely to share a required skill. We cluster POLCA training histories to recover these skill groups. For each datapoint , we compute the total cumulative change in loss along each basis vector by summing over the previous POLCA values. We denote this sum the projected loss .
We obtain 1d projected loss trajectories for breakthrough clustering by computing at every time . As our clustering algorithm, we choose Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) [5] because it distinguishes cluster outliers and discovers clusters with variable density (i.e., similarly shaped curves that lay far apart in their metrics).

### v4

@@ Abstract
Loss curves are smooth during most of model training, so visible discontinuities stand out as possible conceptual breakthroughs. These breakthroughs enable a deeper understanding of the model’s concept structure, but only when they are properly identified. This paper argues that similar breakthroughs occur frequently throughout training, but they are obscured by a loss metric that collapses all variation into a single scalar. To find these hidden transitions, we introduce POLCA, a method for decomposing changes in loss along arbitrary bases of the low-rank training subspace. We use our method to identify clusters of samples that share similar changes in loss during training, disaggregating the overall loss into that of smaller groups of conceptually similar data. We validate our method on synthetic arithmetic and English language modeling, showing that POLCA recovers clusters that represent interpretable breakthroughs in the model’s capabilities. We demonstrate the promise of these hidden breakthroughs as a tool for unsupervised interpretability.
@@ 1 Introduction
As large language models train, various internal structures develop during abrupt breakthroughs. These sudden drops in loss reveal the formation of mechanisms for in-context learning (Olsson et al., 2022b), natural language grammar (Chen et al., 2024a), hierarchical generalization (Murty et al., 2023), and many other concepts (McGrath et al., 2022; Lovering et al., 2022; Power et al., 2022; Abbe et al., 2021). However, the loss curve as a whole remains stubbornly smooth. As a result, these momentary conceptual breakthroughs are treated as isolated curiosities, while the majority of training behavior is considered predictable.
These breakthroughs—often colloquially termed phase transitions (Olsson et al., 2022a; Chen et al., 2024a; Murty et al., 2023)—are extremely consequential for our understanding of neural networks. Phase transitions represent critical periods of learning, so they offer key insights for training and optimization. For instance, introducing noisy data or changing the optimizer during a phase transition can significantly reduce the downstream performance of a model (Achille et al., 2017; Chen et al., 2024a). Their timing is used as evidence the role of specific mechanisms in model capabilities (Zhong et al., 2023; Olsson et al., 2022b; Chen et al., 2024a). If we could find more of these momentary training events, it could expand our understanding even further.
Prior work identifies breakthroughs through a top-down approach by measuring the training dynamics of a predefined concept or skill and searching for sudden changes. We instead propose a bottom-up unsupervised method for finding breakthroughs by grouping data points that have similar training behavior. This data-centric approach can be used to inform optimization choices such as data selection or learning rate scheduling. Like other bottom-up interpretability methods such as SAEs, PCA, and transcoders, our method seeks concepts that are used naturally by the model, rather than imposing an assumed structure onto learning and representation.
This work shows that in fact, a model undergoes many breakthroughs during training, but most are concealed when averaging all data into a single loss curve. Instead of averaging, we divide up the loss curve in two different ways to find hidden breakthroughs. First, we disaggregate the aggregate loss into losses on individual examples. By clustering the individual loss curves, we identify subsets of data that experience synchronized changes in loss, implying that they rely on the same conceptual breakthrough. However, any individual example might benefit from multiple breakthroughs; such an example may undergo changes synchronized with different data subsets at different times. Furthermore, distinct concepts might appear simultaneously, erroneously merging their data clusters. To recover the underlying learned concepts, we may have to identify multiple separate breakthroughs which affect the loss curve of a single example.
@@
We show that some learned concepts can be identified by clustering exact loss, while others cannot (Section 4.2). Using POLCA, we extend our cluster analysis to identify these hidden conceptual breakthroughs obscured in the exact loss curves. We automatically identify specific concepts learned during breakthroughs in both synthetic (Section 4) and natural language settings (Section 5).
@@ 2 Background: How much can we learn from learning dynamics?
Various loss breakthroughs have been interpreted as learning specific concepts. But why expect additional interpretable breakthroughs to underlie periods of undifferentiated, gradual model improvement? Our approach is justified by the nature of the loss surface’s complexity, illustrated by Figure 1, in which a smooth curve emerges by eliding breakthroughs from each dimension.
Why expect multiple breakthroughs?
A very early phase transition is to be expected early in training after a brief memorization stage (Shwartz-Ziv & Tishby, 2017). In this sense, the most celebrated breakthroughs—those signaling the formation of induction heads or arithmetic grokking—are not the first drops in their training curves. Some breakthroughs even depend on earlier breakthroughs, as observed in synthetic tasks (Abbe et al., 2021) and in grammar acquisition (Chen et al., 2024a). If one concept depends on another, each must appear at a different timestep, requiring multiple breakthroughs. Furthermore, as shown by Saxe et al. (2019), summing many phase transitions can result in a smooth curve, supporting our hypothesis that these breakthroughs can appear in stable regions of the loss curve.
Multiple breakthroughs can also come from differences in gradient scale along different directions. Ma et al. (2022) even attributed the early edge-of-stability phase transition (Jastrzȩbski et al., 2020; Cohen et al., 2022) to multiscale structure of the loss surface and, furthermore, noted that this multiscale structure emerges at the range where models become singular: their loss lacks a quadratic approximation in terms of model parameters, creating conditions for breakthroughs under Singular Learning Theory (Watanabe, 2010; Wei et al., 2020; Wang et al., 2024). They argued that this structure is the product of both nonuniform data and nonconvex objectives, respectively justifying the disaggregation and decomposition which we apply to interpret training dynamics.
@@
We track learning on training datapoints and subpopulations, rather than the whole training set, because relevant skills can be acquired at different rates (Arora & Goyal, 2023; Chen et al., 2024b). Individual samples thus exhibit changes in loss out of line with the monotonic average trend (Xia et al., 2023; Rosenfeld & Risteski, 2024). In full-batch gradient descent, Cohen et al. (2022) identified non-monotonicity arising from oscillation about the maximum Hessian eigenvector. Rosenfeld & Risteski (2024) demonstrated that these oscillations occur across different axes for different samples and identified the primary cause: surprisingly human-interpretable semantic features. Even when the loss seems stable, performance can oscillate on edge cases until the model develops relevant capabilities (Qin et al., 2024; Bhaskar et al., 2024). We hypothesize that oscillation represents competing skills that are relevant to different subsets of data. To test this hypothesis, and to interpret the meaning of these directions, we disaggregate the loss into clusters with similar dynamics.
@@
Michaud et al. (2024) analyzed the scaling behavior of models with respect to individual tokens and identified a limitation of token-wise analysis of breakthroughs, which they called polygenic scaling effects—samples which combine multiple skills and therefore exhibit breakthroughs at multiple scales. Our POLCA decomposition directly addresses this limitation by decomposing the loss of each token along multiple basis vectors. If we assume that a specific skill is enabled by movement along that skill’s basis vector, then the loss change attributed to movement along that vector will accelerate at the moment the skill is acquired, for every sample that requires that skill. In this manner, the sample transitions from early to late dynamics through a basis-specific loss breakthrough. In other words, by monitoring changes in directions corresponding to specific skills, we support the speculation of Nanda et al. (2023) that “phase transitions are everywhere.”
@@
In practice, a conceptual breakthrough might not occur in a single direction that persists throughout training. However, there is abundant evidence that linear bases of the low-rank (Gur-Ari et al., 2018) training subspace are conceptually meaningful. In the late stages of training, loss is convex on the line connecting a pair of checkpoints (Frankle et al., 2020) if those checkpoints express similar capabilities (Juneja et al., 2023) and mechanisms (Lubana et al., 2023). If a pair of high-dimensional models lack this linear connection, they still connect nonlinearly (Draxler et al., 2019); however, while parameter settings sampled from their linear connections improve broadly on the capabilities of the original models, those sampled from their nonlinear connections are less robust than the originals (Juneja et al., 2023, ref Appendix G). These observations suggest that linear decomposition should preserve meaningful conceptual features on the loss surface, and our experiments show that the resulting directions are interpretable in practice.
@@ 3 Methods
The key to our approach is the separate consideration of each example’s datapoint loss changes throughout training. We contrast this individualized metric with aggregated loss across an entire dataset. Using the datapoint loss, we can cluster individual example on the basis of its loss , change in loss , or magnitude of change during training.
We next decompose the loss itself into specific directions in the weight space, motivated by several considerations: First, while we have moved from an aggregated loss metric to a more granular datapoint loss metric, we are still only considering breakthroughs that are general enough to be perceived in loss curves. Second, an individual datapoint may benefit from a variety of conceptual breakthroughs, but will not be clustered on the breakthroughs individually. Finally, once we have identified a subset of the data as benefiting from a particular conceptual breakthrough, decomposing into individual weight directions allows us to locate where in the weights the breakthrough occurs and to thereby identify the mechanism involved.
When we break the exact loss curve down by directional movement during training, the resulting decomposed loss will reveal breakthroughs that are specific to a given direction. Our procedure follows three steps: (1) select a basis, (2) decompose the loss along that basis to highlight particular learning events, (3) cluster datapoints according to their shared learning events.
@@ 3.1 Finding the basis
To decompose the loss, we first require an interpretable orthogonal basis. We efficiently compute the eigenvectors of the Hessian matrix using CoLA (Potapczynski et al., 2023) to construct a restricted training subspace. We expect this basis to be interpretable because each basis vector captures a large gradient covariance and therefore represents a potential decision boundary. We select this basis for interpretability, but our approach can use an arbitrary choice of basis—for example, one which targets a particular use case or efficiently leverages optimizer preconditioner values.
@@
The basis is constructed as shown in Algorithm 1. Given intermediate training checkpoints and a number of eigenvectors to compute at each checkpoint, we seek a low rank -dimensional subspace which captures most of the movement during optimization (Gur-Ari et al., 2018). We construct this basis iteratively, starting with : at each checkpoint , we take checkpoint weights and project their loss Hessian onto the nullspace of . From the resulting projection, we append the top eigenvectors to . We compute the eigenvectors using Hessian-vector products Golmant et al. (2018) to avoid explicitly constructing the full Hessian matrix. The resulting basis is designed to include directions of highest curvature at each checkpoint so that it will capture synchronized loss behavior throughout training. Note that the very top eigenvectors are likely to reflect local oscillation, rather than conceptually meaningful long-term movement (Song et al., 2024), but as we continue to add to the low rank basis, we include more directions of long-term stable movement. We discard the oscillatory directions which do not provide an overall decrease in loss over the course of training according to POLCA by removing the directions with an increase in the mean projected loss from checkpoint to . In this manner, we first construct a basis based on local information, then filter out directions that do not represent long-term movement. This construction finds local high-curvature directions that may be important for breakthroughs in the intermediate stages of training while ensuring that the basis does not overfit to local oscillations.
3.2 Decomposing the loss with POLCA
@@
To decompose the loss along our basis, we propose a modified version of Loss Change Allocation (LCA) (Lan et al., 2020) which we call Projection-Oriented Loss Change Allocation (POLCA). LCA is a tool for analyzing changes in aggregated loss on dataset between two checkpoints. The output of LCA is the empirical loss change between a pair of checkpoints which can be attributed to the motion of each individual weight unit. Given two consecutive checkpoints with parameters and , LCA reformulates the change in loss as its first-order Taylor approximation, a sum of the loss changes attributed to the movement of each individual model parameter :
@@
The POLCA decomposition differs from LCA in three key ways. First, we do not restrict each direction to correspond to a single unit , instead permitting an arbitrary orthonormal basis vector to replace the axis-aligned basis vectors in LCA; we project onto this basis vector using the dot product . Second, we are interested in changes in the loss on each individual example , not the entire dataset . These first two modifications provide the first-order POLCA decomposition:
@@
The third key difference is that we use a second-order approximation because this basis is constructed explicitly from the Hessian eigenvectors. To understand why this choice of basis warrants a second-order approximation, recall that each basis vector is an eigenvector of the Hessian matrix at some timestep , where is chosen because it has the largest eigenvalue over the whole dataset. If we assume that the top eigenvectors of the aggregate Hessian maintain high curvature at other points in training and on individual datapoints, then the scaling factor in the second-order Taylor term will be very large even at the datapoint level. Limiting the approximation to only the first order term gives poor guarantees on error, as the second-order term could be expected to dominate. Although empirically the difference between the first and second-order values is small (see Appendix I), we nonetheless guarantee a better estimate due to lower Lagrange error bounds by computing the second-order approximation below.
@@
Exact computation of the second-order term would be intractable, requiring computation of the top eigenvalues/vectors for each individual datapoint . Instead, we can approximate it by substituting the true eigenvalue, denoted , with the curvature of the individual loss in the direction , i.e. . If the aggregate Hessian eigenvector is close to the span of the top eigenvectors of the datapoint-specific Hessian for , this provides a reasonable estimate while reducing calculation to a single Hessian-vector product per eigenvector. We therefore approximate the basis projection of the datapoint Hessian as derived in Appendix C.
@@
Equipped with this second-order approximation of the datapoint Hessian’s projection onto our basis, we account for the high curvature and possible domination by the higher order term by modifying Equation 2 into the second-order Taylor expansion using the approximation from Equation 5. We can compute this second-order term with limited additional computational complexity by keeping track of the eigenvalues for each Hessian eigenvector and the aggregate gradient at each checkpoint.
@@ 3.3 Clustering the loss
POLCA, above, provides curves that show how a decomposed loss changes with respect to each training example. We assume that if several examples show similarly timed loss changes in the same direction, they likely rely on the same conceptual breakthroughs or learning events; therefore, they are likely to share a required skill, or specific capability needed for a given task. We cluster POLCA training histories to recover these skill groups. For each datapoint , we compute the total cumulative change in loss along each basis vector by summing over the previous POLCA values. We denote this sum the projected loss .
We obtain 1d projected loss trajectories for breakthrough clustering by computing at every time . We cluster trajectories using Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) Campello et al. (2013) because it distinguishes cluster outliers and discovers clusters with variable density (i.e., similarly shaped curves that lay far apart in their metrics). We cluster the trajectories for each basis vector separately to ensure that the clustering can capture multiple skills per token.
3.4 Defining and identifying hidden breakthroughs
We use POLCA to recover hidden breakthroughs in training, so we must quantify whether clustered trajectories are distinguished by breakthroughs. We use the formulation defined by Chen et al. (2024b) to compute the start of a breakthrough in a given function for a given datapoint and basis vector :
Here, approximates the maximum point of acceleration of in . can be either the projected loss for a given basis vector or the exact loss . We define a hidden breakthrough as a breakthrough that occurs in the flat region of the exact loss curve. That is, if we set a threshold beyond which the exact loss curve is flat, then a given set has a hidden breakthrough in a metric if the expected value of the start of breakthroughs in is greater than :